# Laypeople’s Egocentric Perceptions of Copyright for AI-Generated Art

### v1

@@ Abstract.
Recent breakthroughs in generative AI (GenAI) have fueled debates concerning the status of AI-generated creations under copyright law. This research investigates laypeople’s perceptions ( = 424) of AI-generated art concerning factors associated with copyright protection. Inspired by prior work suggesting that people show egocentric biases when evaluating their own creative outputs, we also test if the same holds for AI-generated art. Namely, we study the differences between the perceptions of those who have something to gain from copyright protection—creators of AI-generated art—and uninvested third parties.
To answer our research questions, we held an incentivized AI art competition, in which some participants used a GenAI model to generate images for consideration while others evaluated these submissions. We find that participants are most likely to attribute authorship and copyright over AI-generated images to the users who prompted the AI system to generate the image and the artists whose creations were used for training the AI model. We also find that participants egocentrically favored their own art over other participants’ art and rated their own creations higher than other people evaluated them. Moreover, our results suggest that people judge their own AI-generated art more favorably with respect to some factors (creativity and effort) but not others (skills). Our findings have implications for future debates concerning the potential copyright protection of AI-generated outputs.
@@ 1. Introduction
Recent breakthroughs in generative artificial intelligence (GenAI) have pushed the boundaries of what machines can generate across various domains. Text-based models (Achiam et al., 2023) have become pervasive online, enabling users to generate long-form texts through short prompts and revolutionizing how one searches for information online. Similarly, AI image models (Betker et al., 2023) have empowered its users to generate highly realistic and detailed images from short textual descriptions. Multimodal GenAI models (Team et al., 2023) show even more promise, generating text, images, audio, and other types of data. These systems learn from large datasets of human creations how to generate text, images, and other content that may be indistinguishable from their human-created counterparts (Köbis and Mossink, 2021; Longoni et al., 2022; Jakesch et al., 2023).
Although GenAI models have the potential to revolutionize how humans express their creativity, they also pose novel challenges to society. One domain that has received considerable attention in the context of GenAI is copyright law. Copyright law regulates works of authorship, such as paintings and novels, and determines who should have exclusive rights over these creations. Extensive literature has examined how copyright law should address AI-generated works (e.g., (Guadamuz, 2017; Gervais, 2019; Smits and Borghuis, 2022; Eshraghian, 2020; Lee et al., 2023), and several lawsuits are currently underway (O’Brien, 2024; Davis, 2023)) to determine whether training on copyrighted material violates the law and if AI-generated content warrants the same legal protection as human-created works.
Discussions surrounding copyright protection of AI-generated outputs have been primarily normative (e.g., (Guadamuz, 2017; Gervais, 2019)), with little focus on capturing the opinion of GenAI users. Scholars debate whether AI-generated outputs are eligible for copyright protection (Smits and Borghuis, 2022) and, if so, who should hold the rights associated with this protection (Eshraghian, 2020). Here, we investigate laypeople’s expectations of copyright law in relation to GenAI outputs.
@@
Finally, empirical studies capturing the opinion of laypeople about copyright is distinctively important in the context of GenAI. GenAI aims to democratize the ability to create works that could be eligible for copyright protection—something that used to be restricted to skilled artists—enabling non-artists to participate in the creative society.
@@
This paper examines laypeople’s perceptions ( = 424) of AI-generated images vis-à-vis their potential copyright protection. We first capture how laypeople judge GenAI images with respect to factors that help determine whether human creations are eligible for copyright in different jurisdictions (Eshraghian, 2020). Furthermore, we investigate whether laypeople believe AI-generated images warrant copyright and, if so, who should own it. More specifically, we address the following research questions:
RQ1)
@@
How do laypeople evaluate AI-generated images concerning the creativity, effort, and skills involved in the creation process?
RQ2)
Who do laypeople believe are the authors of AI-generated images?
RQ3)
@@
Who do laypeople believe should hold the rights to 1) display and 2) make copies of AI-generated images?
@@
To investigate laypeople’s perceptions of AI-generated art, we conducted an experimental study in the form of a juried AI art exhibition. To guide our experimental design, we leveraged prior research that found that people exhibit egocentric biases in their judgments of creative works (Norton et al., 2012; Buccafusco and Sprigman, 2011, 2010), particularly if owned or created by them (Spellman and Schauer, 2008). We hypothesize that similar biases may emerge in people’s perceptions of copyright, since they relate to questions surrounding ownership of creative works.
@@
Participants engaged in the exhibition either as creators (by using an GenAI model to create art), invested evaluators (by generating art and evaluating other people’s submissions), or uninvested evaluators (by only evaluating others’ images). Figure 1 presents a high-level overview of our experimental design. Our between-subjects design allowed us to study how perceptions about copyright vary between those who have something to gain from copyright protection and the exhibition—creators of AI-generated art—and uninvested third parties. For instance, we hypothesized that creators of AI-generated art will egocentrically overestimate the quality of their own creations and exhibit greater support for the exclusive ownership structure that copyright protection could afford them.
@@
Our study was designed to maximize ecological validity. The decision to hold an AI art exhibition was inspired by real-world examples of AI-generated images winning art exhibitions (Roose, 2022) and competitions focused solely on AI art (Anderson, 2014). Our AI art exhibition rewarded the top-10 best submissions, mimicking some of the financial incentives involved in copyright decisions. It also simulates other non-monetary incentives, such as exposure and recognition; selected images were displayed in a website, in which participants had the choice to attach their names to their creations.111https://thegcamilo.github.io/AI-art-exhibition/
@@
Our results suggest that people believe creativity and effort—but not necessarily skills—are necessary to create art using AI. Participants also indicated that users and those whose creations were used to train the GenAI model should be considered authors of AI-generated images and enjoy the rights to display and make copies of them. In contrast, people were less likely to attribute authorship and the rights associated with copyright protection to the AI model itself and the company that developed it.
Focusing on how people’s perceptions vary egocentrically, we found that participants evaluate the process of using GenAI to generate art egocentrically with respect to some factors—e.g., creativity and effort—but not others—such as skills. When deciding who should win the art exhibition’s monetary award, however, participants judged their own creations much more favorably than art generated by others, supporting our egocentric hypotheses. Surprisingly, we found the opposite trend in attributions of authorship to the user of the GenAI model, such that evaluators attributed more authorship to creators than creators themselves. Participants’ opinions concerning who should hold the rights to display and make copies of AI-generated images did not vary by their role in the AI art exhibition.
@@
Our research has implications for the deployment of GenAI models and their future regulation under copyright law. Our findings call for the consideration of a more distributed ownership structure of copyright, under which training data contributors are also recognized as authors and rights-holders. People’s attribution of authorship and rights to data contributors rather than the company that developed the AI model raise questions concerning current business models that concentrate profits in corporate entities at the expense of human artists (Chayka, 2023; Merkley, 2023). We discuss how existing legal doctrines (e.g., neighbouring rights, licensing models) and computer science research could help ensure that training data contributors are compensated.
@@
Our findings suggest that egocentric biases become relevant in perceptions of GenAI outputs and their associated copyright when monetary incentives come into play. Although participants did not prioritize themselves when asked to indicate who should hold hypothetical rights over their creations, they overestimated the quality of their images when that determined real monetary rewards. Our evidence of egocentric biases suggests that some conflicts of interest may arise in discussions surrounding the legal status of AI-generated art under copyright law.
@@ 2. Background
One particular domain that has been directly impacted by the emergence of GenAI is intellectual property (IP) law. IP law refers to the rules that regulate the rights associated with human creations, such as inventions and literary and artistic works, determining who should control and benefit from them. How to deal with AI-generated outputs regarding IP rights remains an open question. For instance, should machine-generated artistic works be protected similarly to their human-created counterparts? Who should enjoy the rights that would normally be associated with this creation?
@@
Debates on how to address novel technologies with copyright law is not a new development. The rise of photography also challenged what was eligible for copyright protection (Hughes, 2011). Similarly, the law was initially unprepared to deal with the emergence of digital art (Geller, 2000) and video-sharing platforms that allowed users to upload copyrighted content without much restriction, leading to solutions that ensure that the creator rights are protected (e.g., Google’s content ID (Keene, 2012)). Now it is the time of GenAI to contest what warrants (or not) copyright protection.
The challenges posed by GenAI to copyright law can be grouped in three overarching questions (Franceschelli and Musolesi, 2022), all of which we discuss below: 1) does training AI models on copyrighted data infringe on the copyright of the training data?; 2) are AI-generated outputs eligible for copyright protection; and if so 3) who owns the copyright? In this paper, we focus on the two latter questions. We also discuss some potential alternative regulatory frameworks for GenAI outputs, as well as prior work on lay perceptions of copyright and GenAI.
2.1. Does Training GenAI Infringe on the Copyright of the Training Data?
In contrast, proponents of GenAI assert that training does not infringe on the copyright of its training data. Such arguments in favor of GenAI often rely on the United States’ fair use doctrine (Lemley and Casey, 2020), which permits limited use of copyrighted materials under specific conditions, and the text and data mining copyright exceptions in the European Union (EU) (Selvadurai and Matulionyte, 2020). The wide availability of GenAI models have triggered several lawsuits that are currently underway (e.g., (O’Brien, 2024)) and will decide whether training GenAI with copyrighted material without compensating its owners is legal under copyright law.
@@ 2.2. Can AI-Generated Outputs Be Protected by Copyright?
Another question raised by GenAI concerning copyright law refers to whether AI-generated outputs should receive the same protection that are assigned to works that are created solely by humans. If works generated with the assistance of an AI model are eligible for copyright protection, someone would have exclusive rights over it; on the other hand, if they are not eligible, these outputs would be part of the public domain, meaning that anyone would be able to use these works without permission.
Scholars largely disagree on this particular question. Some argue extending copyright law to machine-created works would reduce the value of human creativity (Merkley, 2023), flood the market with creations of questionable quality (Gervais, 2019), and concentrate power in the hands of a few (Iaia, 2022; Chan et al., 2023). In contrast, proponents of extending copyright to AI-generated works defend that protecting these outputs could promote innovation by incentivising research and development of AI (Hristov, 2016; Kop, 2019), as well as enable users to create works that would not be possible without GenAI.
@@ 2.3. Who Owns the Potential Copyright of AI-Generated Outputs?
If an AI-generated work is granted copyright, some legal entity would have exclusive rights over it; but who would that be? A GenAI output is the result of a collaboration between several actors, including the model’s developers, its users, those who potentially own the training data, and the AI model itself, making it difficult to determine who should own it.333This question is an instance of the “problem of many hands” (Van de Poel, 2015), which posits that it is difficult to determine who is ultimately responsible for collective actions. Scholars have also explored how AI may complicate this search for a responsible actor, particularly when it causes harm (Coeckelbergh, 2020; Matthias, 2004).
A common proposition is that the user of a GenAI model, i.e., the person who gave it instructions, should be granted exclusive rights over its outputs (Eshraghian, 2020). However, it is not clear whether this user would satisfy the conditions that determine authorship and, hence, copyright ownership. The US Copyright Office has stated that merely prompting an AI model does not qualify the user for authorship; instead, it proposes a case-by-case analysis that would determine whether the work contains “sufficient human authorship” (e.g., whether the user “select[ed] or arrange[d] AI-generated material in a sufficiently creative way” (Office, 2023)). Another possibility would be granting copyright to the “the person by whom the arrangements necessary for the creation of the work are undertaken,” as proposed by the UK Intellectual Property Office (UK Intellectual Property Office, 2022), which could interpreted as the developer of the AI model (Ihalainen, 2018). Other proposals include considering the AI model itself as an author (and thus copyright holder) (Smits and Borghuis, 2022; Lima et al., 2020) or granting a form of joint authorship (i.e., collective ownership) to the many entities involved (Khosrowi et al., 2023). This research investigates who laypeople consider to be authors of AI-generated images (RQ2), as well as who they believe should have the rights to display and make copies of them (RQ3).
2.4. Alternatives to Copyright Law
Although copyright law often takes central stage in the discussion surrounding ownership of GenAI outputs, it is not exempt from critiques. Copyright law is often criticized by its potential to hinder innovation by concentrating power in monopolies (Ihalainen, 2018; Levendowski, 2018). Furthermore, extending copyright law to GenAI could conflict with its primary objective—to promote human creativity; if AI-generated works are protected, they could devalue human creativity by flooding the market with artificial competition to human creations (Iaia, 2022).
Scholars have proposed different approaches for AI-generated works to promote innovation. For instance, the reproduction of GenAI models and their outputs could be restricted under a suis generis doctrine, which grants exclusive rights in situations in which there has been substantial financial investment (Smits and Borghuis, 2022). Someone other than the author—which is hard to define in the case of AI-generative works, as discussed above—could also have “neighbouring” rights over AI outputs as a way to protect their investment (Lee, 2021; Iaia, 2022). AI-generated outputs could also be covered by other branches of IP law, such as trademark law, which relies on owners maintaining and enforcing their rights (Ihalainen, 2018). Another approach could be moving away from copyright law’s focus on creativity and originality as grounds for ownership, and instead reward actors who put effort and demonstrate skill when using GenAI (Guadamuz, 2017).

### v2

# Public Opinions About Copyright for AI-Generated Art: The Role of Egocentricity, Competition, and Experience

@@ Abstract.
Breakthroughs in generative AI (GenAI) have fueled debates concerning the artistic and legal status of AI-generated creations. We investigate laypeople’s perceptions () of AI-generated art through the lens of copyright law. We study lay judgments of GenAI images concerning several copyright-related factors and capture people’s opinions of who should be the authors and rights-holders of AI-generated images. To do so, we held an incentivized AI art competition in which some participants used a GenAI model to create art while others evaluated these images. We find that participants believe creativity and effort, but not skills, are needed to create AI-generated art. Participants were most likely to attribute authorship and copyright to the AI model’s users and to the artists whose creations were used for training. We find evidence of egocentric effects: participants favored their own art with respect to quality, creativity, and effort—particularly when these assessments determined real monetary awards.
@@ 1. Introduction
Recent breakthroughs in generative artificial intelligence (GenAI) have pushed the boundaries of what machines can generate across various domains. Text-based models (Achiam et al., 2023) have become pervasive online, enabling users to generate long-form texts from short prompts and revolutionizing how one searches for information online. Similarly, AI image models (Betker et al., 2023) have empowered their users to generate highly realistic and detailed images from short textual descriptions. Multimodal GenAI models (Team et al., 2023) show even more promise, generating text, images, audio, and other types of data. These systems learn from large datasets of human creations how to generate text, images, and other content that may be indistinguishable from their human-created counterparts (Köbis and Mossink, 2021; Longoni et al., 2022; Jakesch et al., 2023). GenAI models have now become part of several real-world products (Fraser, 2024), such as search engines (Reid, 2024), online shopping websites (Amazon, 2024), and messaging apps (WhatsAp, 2024).
Although GenAI models have the potential to revolutionize how humans interact with technology and express their creativity, they also pose novel challenges to society. For instance, there exist debates concerning whether GenAI outputs could devalue human creativity (Merkley, 2023), fuel artistic appropriation (Goetze, 2024), and concentrate power in the hands of a few corporate actors (Iaia, 2022; Chan et al., 2023). Such debates are both philosophical—concerning aesthetics and societal perceptions of what art is and is not—and legal. Copyright law, the body of law that regulates works of authorship (e.g., paintings and novels) and determines who should have exclusive rights over these creations, has received considerable attention in the context of GenAI. Extensive literature has examined how copyright law should address AI-generated works (e.g., (Guadamuz, 2017; Gervais, 2019; Smits and Borghuis, 2022; Eshraghian, 2020; Lee et al., 2023), and several lawsuits are currently underway (O’Brien, 2024; Davis, 2023) to determine whether training on copyrighted material violates the law and if AI-generated content warrants the same legal protection as human-created works.
Thus far, considerations of AI-generated outputs from a copyright lens have been primarily normative (e.g., (Guadamuz, 2017; Gervais, 2019)), with little focus on capturing the opinion of GenAI users. Scholars debate whether AI-generated outputs are eligible for copyright protection (Smits and Borghuis, 2022) and, if so, who should hold the rights associated with this protection (Eshraghian, 2020). Here, we investigate laypeople’s perceptions of GenAI art through a copyright lens.
The figure is a flowchart depicting the three experimental conditions: creator, invested evaluator, and uninvested evaluator. Creators generate an image for consideration and then evaluate their own creation. Invested evaluators generate an image, evaluate their own creation, and then evaluate images generated by creators. Uninvested evaluators only evaluate images generated by creators.
@@
Third, empirical studies capturing the opinion of laypeople about copyright are distinctively important in the context of GenAI. GenAI aims to democratize the ability to create works that could be eligible for copyright protection—something that used to be restricted to skilled artists—enabling non-artists to participate in the creative society.
Research on how novel digital systems impact people’s perceptions of ownership and shape individuals’ processes for meaning-making around the digital artifacts produced by those same systems is core to human-computer interaction (HCI) (Xu and Warschauer, 2020; Gulotta et al., 2013). We take the law as a perspective lens through which we explore people’s perceptions of one of the digital artifacts of GenAI systems: AI-generated art. Because regulations constraint what users and HCI designers can do, capturing people’s perceptions of technology through the lens of the law sheds light on how expectations might clash with policy (Awad et al., 2020) and provides implications for designers aiming to develop systems that function within legal constraints while satisfying user needs (Fiesler, 2020). In view of the recent developments in GenAI and how this technology has become embedded in both HCI research and real-world products, our research aims to inform the HCI community towards the design of GenAI systems that are both lawful and fulfilling to users (Lazar et al., 2019).
We acknowledge that debates on how novel technologies can change the meaning of existing cultural artifacts and their regulation are not a new development. The rise of photography also challenged what was eligible for copyright protection (Hughes, 2011). Similarly, the law was initially unprepared to deal with the emergence of digital art (Geller, 2000) and video-sharing platforms that allowed users to upload copyrighted content without much restriction, leading to solutions that ensure that the creator rights are protected (e.g., Google’s content ID (Keene, 2012)). Now, GenAI poses novel policy questions by contesting what warrants (or not) copyright protection—the topic of the research we present below.
1.1. Empirical Study of AI-Generated Art Through the Lens of Copyright
@@
This paper examines laypeople’s perceptions ( = 432) of AI-generated images in relation to their potential copyright protection. We first capture how laypeople judge GenAI images with respect to factors that help determine whether human creations are eligible for copyright in different jurisdictions (Eshraghian, 2020) (see Section 2.1 for more detailed legal background). Then, we investigate whether laypeople believe AI-generated images warrant copyright and, if so, who should own them (see Section 2.2 for more detailed legal background). More specifically, we address the following research questions:
@@
RQ1: How do laypeople evaluate AI-generated images concerning the creativity, effort, and skills involved in the creation process?
RQ2: Who do laypeople believe are the authors of AI-generated images?
@@
RQ3: Who do laypeople believe should hold the rights to 1) display and 2) make copies of AI-generated images?
@@
To investigate laypeople’s perceptions of AI-generated art, we conducted an experimental study in the form of a juried AI art exhibition. To guide our experimental design, we leveraged prior research that found that people judge their own creations (or the creations they own) differently than the creations of others—the literature often terms this “egocentric bias” (Norton et al., 2012; Buccafusco and Sprigman, 2011, 2010; Spellman and Schauer, 2008). We hypothesize that similar biases may emerge in people’s perceptions of copyright since they relate to questions surrounding ownership of creative works.
@@
Participants engaged in the exhibition either as creators (by using a GenAI model to create art), invested evaluators (by generating art and evaluating other people’s submissions), or uninvested evaluators (by only evaluating others’ images). Figure 1 presents a high-level overview of our experimental design. Our between-subjects design allowed us to study how perceptions about copyright vary between those who have something to gain from copyright protection and the exhibition—creators of AI-generated art—and uninvested third parties. For instance, we hypothesized that creators of AI-generated art would egocentrically overestimate the quality of their own creations and exhibit greater support for the exclusive ownership structure that copyright protection could afford them.
@@
Our study was designed to maximize ecological validity. The decision to hold an AI art exhibition was inspired by real-world examples of AI-generated images winning art exhibitions (Roose, 2022) and competitions focused solely on AI art (Anderson, 2014). Our AI art exhibition rewarded the top-10 best submissions, mimicking some of the financial incentives involved in copyright decisions. It also simulates other non-monetary incentives, such as exposure and recognition; selected images were displayed on a website, in which participants had the choice to attach their names to their creations.111https://thegcamilo.github.io/AI-art-exhibition/
1.2. Findings & Implications
@@
Our results suggest that people believe creativity and effort—but not necessarily skill—are necessary to create art using AI (RQ1). Participants also indicated that users and those whose creations were used to train the GenAI model should be considered authors of AI-generated images (RQ2) and enjoy the rights to display and make copies of them (RQ3). In contrast, people were less likely to attribute authorship (RQ2) and the rights associated with copyright protection (RQ3) to the AI model itself and the company that developed it.
@@
Our research has implications for the development of GenAI models and their future regulation under copyright law. Our findings call for the consideration of a more distributed ownership structure of copyright, under which training data contributors are also recognized as authors and rights-holders. People’s attribution of authorship and rights to data contributors rather than the company that developed the AI model raises questions concerning current business models that concentrate profits in corporate entities at the expense of human artists (Chayka, 2023; Merkley, 2023). Our findings interrogate current practices of GenAI model designers, who may—knowingly or accidentally—promote the interests of powerful corporate actors to the detriment of artists’ welfare. We discuss how existing legal doctrines (e.g., neighbouring rights, licensing models), computer science research, and self-questioning from those who design GenAI products could help ensure that training data contributors are compensated.
@@
Focusing on how people’s perceptions vary egocentrically, we found that participants evaluate the process of using GenAI to generate art egocentrically with respect to some factors—e.g., creativity and effort—but not others—such as skills. Our results indicate that egocentric biases become particularly relevant when monetary incentives come into play. Although participants did not prioritize themselves when asked to indicate who should hold hypothetical rights over their creations, they overestimated the quality of their images when that determined real monetary rewards. Namely, when deciding who should win the art exhibition’s monetary award, participants judged their own creations much more favorably than art generated by others, supporting our egocentric hypotheses. Surprisingly, we found the opposite trend in attributions of authorship to the user of the GenAI model. Creators were less likely to rate themselves as authors than evaluators, potentially suggesting that GenAI users are aware that authorship can bring both benefits and liability for their creations. Our evidence of egocentric biases suggests that some conflicts of interest may arise in discussions surrounding the legal status of AI-generated art under copyright law. We also discuss how decisions concerning the terms of use of real-world GenAI systems may shape people’s expectations of AI-generated images in light of their potential financial returns.
@@ 2. Background
One particular domain that has been directly impacted by the emergence of GenAI is intellectual property (IP) law. IP law refers to the rules that regulate the rights associated with human creations, such as inventions and literary and artistic works, determining who should control and benefit from them. How to deal with AI-generated outputs regarding IP rights remains an open question. For instance, should machine-generated artistic works be protected similarly to their human-created counterparts? Who should enjoy the rights that would normally be associated with these creations?
@@
The challenges posed by GenAI to copyright law can be grouped in three overarching questions (Franceschelli and Musolesi, 2022): 1) does training AI models on copyrighted data infringe on the copyright of the training data?; 2) are AI-generated outputs eligible for copyright protection; and if so 3) who owns the copyright? In this paper, we focus on the two latter questions, both of which we discuss in §2.1-2.2 below.333We also present the debate on how training GenAI models might violate the copyright of the training data (i.e., question #1 above), as well as some potential alternative regulatory frameworks for GenAI in the Appendix for conciseness. Furthermore, we present prior work on lay perceptions of copyright (§2.3) and GenAI (§2.4) and motivate our focus on potential biases that may emerge in the context of copyright and GenAI (§2.5).
@@ 2.1. Can AI-Generated Outputs Be Protected by Copyright?
An important question raised by GenAI concerning copyright law refers to whether AI-generated outputs should receive the same protection that is assigned to works that are created solely by humans. If works generated with the assistance of an AI model are eligible for copyright protection, someone would have exclusive rights over it; on the other hand, if they are not eligible, these outputs would be part of the public domain, meaning that anyone would be able to use these works without permission.
Scholars disagree on this particular question. Some argue extending copyright law to machine-created works would reduce the value of human creativity (Merkley, 2023), flood the market with creations of questionable quality (Gervais, 2019), and concentrate power in the hands of a few (Iaia, 2022; Chan et al., 2023). In contrast, proponents of extending copyright to AI-generated works defend that protecting these outputs could promote innovation by incentivizing research and development of AI (Hristov, 2016; Kop, 2019), as well as enable users to create works that would not be possible without GenAI.
@@ 2.2. Who Owns the Potential Copyright of AI-Generated Outputs?
If an AI-generated work is granted copyright, some legal entity would have exclusive rights over it; but who would that be? A GenAI output is the result of a collaboration between several actors, including the model’s developers, its users, those who potentially own the training data, and the AI model itself, making it difficult to determine who should own it.444This question is an instance of the “problem of many hands” (Van de Poel, 2015), which posits that it is difficult to determine who is ultimately responsible for collective actions. Scholars have also explored how AI may complicate this search for a responsible actor, particularly when it causes harm (Coeckelbergh, 2020; Matthias, 2004).
A common proposition is that the user of a GenAI model, i.e., the person who gave it instructions, should be granted exclusive rights over its outputs (Eshraghian, 2020). However, it is not clear whether this user would satisfy the conditions that determine authorship and, hence, copyright ownership. The US Copyright Office has stated that merely prompting an AI model does not qualify the user for authorship; instead, it proposes a case-by-case analysis that would determine whether the work contains “sufficient human authorship” (e.g., whether the user “select[ed] or arrange[d] AI-generated material in a sufficiently creative way” (Office, 2023)). Another possibility would be granting copyright to the “the person by whom the arrangements necessary for the creation of the work are undertaken,” as proposed by the UK Intellectual Property Office (UK Intellectual Property Office, 2022), which could interpreted as the developer of the AI model (Ihalainen, 2018). Other proposals include considering the AI model itself as an author (and thus copyright holder) (Smits and Borghuis, 2022; Lima et al., 2020) or granting a form of joint authorship (i.e., collective ownership) to the many entities involved (Khosrowi et al., 2023). This research investigates whom laypeople consider to be authors of AI-generated images (RQ2), as well as who they believe should have the rights to display and make copies of these images (RQ3).